# Hallucination to Truth: A Review of Fact-Checking and Factuality Evaluation in Large Language Models

### v1

@@ 4.3  Datasets for Training and Evaluating Fact-Checking Systems (RQ3)
Benchmark Datasets for RAG-Based Fact Verification. Core claim verification datasets such as FEVER [86], FEVEROUS [9], and HOVER [87] are widely used to evaluate RAG pipelines, where the model is used to retrieve relevant evidence from Wikipedia or structured sources and then generate a verdict. These datasets provide gold-standard evidence, making them ideal for training retrievers and verifying generation accuracy. LIAR [88] and RAWFC [89] further allow the assessment of RAG-based models on political and news-based claims with distinctive complexity and source structures.
Domain-specific Datasets. In domain-specific applications, datasets such as SciFact [90], COVID-Fact [91], MedMCQA [92], BioASQ [93], and PubMedQA [94] are frequently employed in RAG frameworks that are aligned for the biomedical domain. These allow models to retrieve evidence from the medical literature (e.g., via MedRAG) and cross-validate LLM outputs. Given the importance of factual accuracy in these domains, these datasets also serve as valuable benchmarks for evaluating and refining hallucination reduction techniques in sensitive contexts.
Multimodal Datasets. Multimodal datasets such as MM-FEVER [95], Post-4V [96], NewsCLIPpings [74], and MOCHEG [97] allow extended fact-checking to vision-language settings. These are especially relevant for evaluating multimodal RAG systems that allow the use of visual and textual information to assess the veracity of claims. Mismatches between image-caption pairs in these datasets test the models’ ability to detect hallucinated or manipulated content across modalities.
@@
Synthetic and Multilingual Datasets. Synthetic and weak supervision datasets such as ClaimMatch [103], LLM-AGGREFACT [104], and CheckThat22 [105] allow for scalable training and evaluation in low-resource settings. These datasets are often used to pre-train or fine-tune retrievers and scorers within RAG systems, or to assess robustness against adversarial claims and misinformation edits. Additionally, multilingual datasets and some databases, like X-Fact [106], Data Commons Multilingual, and FactStore, help build fact-checking systems that work across different languages. They test whether these systems can find the correct information and give accurate answers, even in non-English settings. This helps make fact-checking tools more significant in global use. Figure 8 visualizes an overview of major dataset types and their domains.
Figure 8: Illustration of major dataset types and domains.
To provide a clear and comprehensive overview, Table 4 lists 72 key datasets used in fact-checking research. It shows which datasets were used for fact checking by using RAG, for hallucination reduction, and their type. Out of 72 datasets, 63 use RAG and 48 are used for hallucination reduction.
Table 4: Datasets and their use in RAG and hallucinations (Halluc.) reduction.
@@ 4.4  Prompt Design, Fine-Tuning, and Domain-Specific Training (RQ4)
Figure 9: Breakdown of approaches in prompt design, fine-tuning, and domain-specific training.
@@ 4.4.1  Basic Prompting Strategies
These strategies primarily rely on internal knowledge and involve presenting the claim or task to the LLMs with minimal or no external context beyond the prompt itself. Their effectiveness is heavily reliant on the model’s pre-trained knowledge, which can be a significant limitation due to the potential for hallucination and outdated information [30, 12, 21].
@@
Zero-shot prompting involves providing the LLMs with only the task description and the input claim, without any specific examples. This can include asking the model to predict the veracity label of a claim directly [12, 72]. Frameworks such as the FACT-AUDIT [45] use zero-shot inference to evaluate the fact-checking capacity of various LLMs. They can yield the lowest average accuracy scores compared to other methods, particularly when relying solely on the model’s internal knowledge. While a simple self-consistency and zero-shot prompt combination was found to be the most effective overall strategy in a multilingual fact-checking study, this effectiveness was strongly tied to the self-consistency decoding strategy, not necessarily the zero-shot nature itself [32]. On the other hand, frameworks such as PCAR [14], which leverage explicit claim decomposition and a dynamic planning mechanism, achieve strong performance in zero-shot settings and outperform other LLM-based approaches, including few-shot and fine-tuned methods. However, reliance solely on internal knowledge for fact-checking is considered unreliable and insufficient, as LLMs are prone to hallucination [30, 12, 21]. Zero-shot prompting without external access means the model must rely on potentially inaccurate or outdated information stored during training [48]. The analysis of rationales generated through zero-shot CoT indicates unreliability for factuality analysis based on internal memorization [35]. The most significant mitigation discussed is incorporating external knowledge retrieval [48, 30, 20, 15, 53, 21].
The Few-Shot Prompting or In-Context Learning (ICL) method involves providing the model with a limited number of examples of the task before presenting the claim to be verified [14, 12]. It leverages the LLM’s ability to learn from examples provided directly in the prompt ("in-context learning") [20]. Few-shot demonstrations are used in methods like HiSS(Hierarchical Step-by-Step) [21] and BiDeV [37] to guide the LLM through multi-step processes. Few-shot-CoT includes example pairs to guide the reasoning process [47, 72].
@@
ICL has been shown to enhance the performance of open-source MLLMs in detecting misinformation, sometimes providing greater improvement than prompt ensembles [57]. Combining ICL with RAG has been shown to improve accuracy in fact verification [53]. However, some sophisticated few-shot ICL methods like Standard Prompting, Vanilla CoT, and ReAct were surpassed by the HiSS method in news claim verification, highlighting the importance of the specific method prompted [48]. While ICL improves performance, open-source MLLMs using ICL still significantly lag behind state-of-the-art proprietary models like GPT-4V. The effectiveness can vary between models and datasets [57]. Manual prompt design for few-shot examples can be heuristic [20]. However, providing examples does not guarantee overcoming reliance on internal knowledge if external information is not integrated [21]. Integrating ICL with RAG or structured, step-by-step prompting frameworks is a key mitigation [53, 21].
@@
Other methods, such as Chain-of-Thought (CoT) prompting, instruct the LLM to output a sequence of intermediate reasoning steps before arriving at the final answer or veracity label [21, 12]. Zero-shot CoT prompts often include eliciting sentences like "Let us think step by step" [12, 32, 47, 16]. This encourages explicit reasoning. Variants of CoT prompting include English CoT (EN-CoT) [32], which focuses on monolingual reasoning, and CoTVP (Chain of Thought Veracity Prediction) [22], which evaluates the truthfulness of the reasoning steps. However, studies show that techniques like CoT, designed to improve reasoning, do not necessarily improve the fact-checking abilities of LLMs. They can even have minimal or negative effects on success rates [32]. Models prompted with CoT may tend to align with the input text rather than verifying its factualness, especially for complex paragraphs, when relying solely on pre-trained knowledge [20]. Vanilla CoT suffers substantially from issues of fact hallucination and omission of necessary thoughts in the reasoning process [21]. Using the LLM for factuality analysis based on its internal memorization, even with zero-shot CoT, indicates unreliability, likely caused by hallucination [12].
@@
The internal mechanism of LLMs to integrate rationales from diverse perspectives via CoT can be ineffective for fake news detection [12]. Integrating CoT with external knowledge retrieval, as in Search-Augmented CoT or ReAct [21], is a key approach to mitigate hallucination and thought omission [21]. Additionally, frameworks that explicitly guide the decomposition and reasoning process, such as FactAgent, are seen as superior to CoT, which primarily acts as a prompting technique [72].
@@
When relying solely on internal knowledge, LLMs prompted with zero-shot, few-shot, or vanilla CoT struggle with fact-checking complex claims and exhibit hallucination [30, 12, 20, 21]. However, accuracy can be low, and improvements in reasoning via CoT alone do not guarantee better fact-checking performance [32]. Techniques forcing binary "true" or "false" judgments also do not enhance overall accuracy [73].
@@ 4.4.2  Prompting Strategies with Integrated External Retrieval
A CoT variant that interleaves reasoning traces with task-specific actions, like querying Google Search or the Wikipedia API, allowing the LLM agent to decide whether to search or continue reasoning based on environmental observations [48, 21, 20, 22]. For instance, by accessing external knowledge, ReAct effectively mitigates hallucination failures compared to vanilla CoT and justifies its reasoning with retrieved citations, enhancing verifiability and explainability. Its performance is highly sensitive to the quality and relevance of search results, and relying solely on internal knowledge when external search fails remains a key limitation [21]. Combining ReAct’s action capabilities with more structured decomposition and step-by-step verification methods, such as HiSS or SELF-CHECKER, can address thought omissions and improve overall performance [21, 20].
@@
Search-Augmented CoT augments vanilla CoT by using the original claim as a search query to retrieve background information, which the LLM incorporates into its thought chain. This approach improves over vanilla CoT by leveraging external knowledge, but can fall short of methods like Standard Prompting or HiSS, which indicates that querying solely with the claim may yield insufficiently detailed results. To mitigate this, more sophisticated query generation strategies and integration methods are needed [21]. HiSS is a few-shot method that prompts the LLM to perform claim verification in fine-grained steps by decomposing claims into subclaims and verifying each step-by-step, raising questions and optionally using web search when confidence is low. It significantly surpasses few-shot ICL counterparts like Standard Prompting, Vanilla CoT, and ReAct in average F1-score, offering superior explainability through enhanced coverage and readability while substantially reducing hallucination and thought omission [21]. However, HiSS still struggles with integrating updated information from mixed sources and incurs high computational costs due to multiple LLM calls, with its performance remaining sensitive to prompt design [20].
@@
Frameworks like SELF-CHECKER [20], BiDeV [37], RAGAR [22], and PACAR [14] integrate prompting and retrieval-augmented generation (RAG) by decomposing fact-checking into subtasks (e.g., claim detection, retrieval, sentence selection, verdict prediction) and using prompts often with few-shot examples to generate search queries, select evidence, and perform step-by-step verification while explicitly incorporating retrieved documents. Incorporating external knowledge via RAG significantly boosts accuracy over internal-only approaches: SELF-CHECKER [20], BiDeV [37], RAGAR [22], and PACAR [14] all show substantial gains, with specialized modules like the Claim Atomizer and Fact-Check-Then-RAG further enhancing predictive power and explanation quality.
@@
RAG-based methods can be hampered by overwhelming context windows [48], outdated or variable search results [20, 22], high computational costs, prompt sensitivity, and manual prompt design [20], and they may still miss fine-grained details even when grounded [30]. To mitigate these issues, current strategies include using IR functions (e.g., BM25) to distill relevant content [30, 48], using multi-agent [76, 37], explicit decomposition and filtering [14, 42, 37], feedback loops [51], and refined RAG processes [30, 42]. Reliance solely on an LLM’s internal knowledge, whether via zero-shot, few-shot, or vanilla CoT prompting, is unreliable for fact-checking and prone to substantial hallucination [21, 30, 12], as zero-shot CoT and vanilla CoT often succumb to memorization pitfalls [12, 16]. Mitigating hallucination primarily involves incorporating external knowledge through ReAct, Search-Augmented CoT, HiSS, and other RAG frameworks [21, 30, 53].
@@
In conclusion, while basic prompting strategies like zero-shot, few-shot, and vanilla CoT offer foundational ways to interact with LLMs for fact-checking, their accuracy and reliability are severely limited by reliance on potentially flawed internal knowledge [30, 12, 21]. The most effective approaches, as highlighted by the sources, involve prompting strategies that explicitly integrate external knowledge retrieval through frameworks like ReAct, HiSS, SELF-CHECKER, BiDeV, RAGAR, and PACAR [21, 48, 20, 53, 37]. These methods use prompting to guide the LLM through processes involving external data access, significantly improving accuracy and mitigating hallucination by grounding the model’s responses in evidence [21, 20, 53].

### v2

@@ 4.3  Datasets for Training and Evaluating Fact-Checking Systems (RQ3)
Benchmark Datasets for RAG-Based Fact Verification. Core claim verification datasets such as FEVER [96], FEVEROUS [9], and HOVER [97] are widely used to evaluate RAG pipelines, where the model is used to retrieve relevant evidence from Wikipedia or structured sources and then generate a verdict. These datasets provide gold-standard evidence, making them ideal for training retrievers and verifying the accuracy of generated outputs. LIAR [98] and RAWFC [99] further allow the assessment of RAG-based models on political and news-based claims with distinctive complexity and source structures.
Domain-specific Datasets. In domain-specific applications, datasets such as SciFact [100], COVID-Fact [101], MedMCQA [102], BioASQ [103], and PubMedQA [104] are frequently employed in RAG frameworks that are aligned with the biomedical domain. These allow models to retrieve evidence from the medical literature (e.g., via MedRAG) and cross-validate LLM outputs. Given the importance of factual accuracy in these domains, these datasets also serve as valuable benchmarks for evaluating and refining hallucination reduction techniques in sensitive contexts.
Multimodal Datasets. Multimodal datasets such as Multimodal-FEVER (Factify) [105], Post-4V [57], NewsCLIPpings [73], and MOCHEG [106] allow extended fact-checking to vision-language settings. These are especially relevant for evaluating multimodal RAG systems that allow the use of visual and textual information to assess the veracity of claims. Mismatches between image-caption pairs in these datasets test the models’ ability to detect hallucinated or manipulated content across modalities.
@@
Synthetic and Multilingual Datasets. Synthetic and weak supervision datasets such as ClaimMatch [112], LLM-AGGREFACT [10], and CheckThat [113] allow for scalable training and evaluation in low-resource settings. These datasets are often used to pre-train or fine-tune retrievers and scorers within RAG systems, or to assess robustness against adversarial claims and misinformation edits. Additionally, multilingual datasets and some databases, like X-Fact [114], Data Commons Multilingual, and FactStore, help build fact-checking systems that work across different languages. They test whether these systems can find the correct information and give accurate answers, even in non-English settings. This helps make fact-checking tools more significant in global use. Figure 9 visualizes an overview of major dataset types and their domains.
Figure 9: Illustration of major dataset types and domains. The bar chart shows the distribution of datasets across different types and application domains, highlighting their relative prevalence and focus areas.
To provide a clear and comprehensive overview, Table 5 lists 72 key datasets used in fact-checking research. It shows which datasets were used for fact-checking by using RAG, for hallucination reduction, and their type. Out of 72 datasets, 63 use RAG, and 48 are used for hallucination reduction. The use cases of each dataset used for fact-checking are grouped into several key types. To test interactive evidence retrieval and conversational-based fact-checking in multi-turn RAG pipelines, dialogue datasets are used [54, 63]. QA datasets test a system’s ability to accurately answer questions by reasoning through open-domain and multi-hop scenarios, where it retrieves and connects relevant information [62, 97]. Fact-checking datasets provide labelled claims with the most reliable evidence for training and benchmarking claim verification [9, 112]. Complex fact-checking in biomedical contexts is typically facilitated by medical and domain-specific datasets (e.g., SciFact, PubMedQA) [63, 104]. LLM and hallucination datasets capture ungrounded outputs to assess truthfulness and hallucination mitigation. Structured data-to-text datasets can produce more accurate texts from tables or knowledge graphs [8]. Summarisation datasets evaluate the production of concise and reliable text [3]. Misinformation datasets help to find and disprove false or manipulated claims [92, 113], and multimodal datasets check that text–image evidence pairs are consistent [43, 90].
Table 5: Datasets and their use in RAG and hallucinations (Halluc.) reduction. It indicates the type of dataset and the tasks for which it has been evaluated.
@@
Table 6: Identified Limitations of Datasets Commonly Used in LLM-Based Fact-Checking Across Different Domains
Dataset Type Key Limitations
Dialogue Low domain diversity and simple claim structures with weak coverage of real misinformation.
Fact-Checking Biased toward Wikipedia claims with poor temporal and geographic variety and slow tracking of emerging misinformation.
QA Encourages benchmark overfitting and lacks adversarial and multi-hop claims.
Medical Expensive annotation with narrow coverage and poor multilingual and low-resource support.
Misinformation / Political Small size and region specific, with rare updates for new narratives.
Structured Data Clean and synthetic, but fails to capture noisy and conflicting real-world data.
Multimodal Sparse and noisy, with weak standards and limited cross-modal reasoning tests.
LLM Mostly synthetic with short and simple prompts and weak domain-specific coverage.
Summarization Subjective references that reward lexical overlap over factual grounding.
Multilingual Uneven language coverage with high-resource bias and weak cultural generalization.
Table 6 represents the key limitations of each dataset type used in fact-checking with LLM. Most datasets are narrow in scope or biased, which affects the reliability of fact-checking in specific domains and languages.
@@ 4.4  Prompt Design, Fine-Tuning, and Domain-Specific Training (RQ4)
Figure 10: Breakdown of approaches in prompt design, fine-tuning, and domain-specific training for fact-checking with LLMs, categorized into four groups. Each group highlights the number of papers reviewed and the representative techniques.
@@ 4.4.1  Basic Prompting Strategies
Basic prompt strategies primarily rely on internal knowledge and involve presenting the claim or task to LLMs with minimal or no external context beyond the prompt itself. Their effectiveness is heavily dependent on the pre-trained knowledge of the model, which can be a significant limitation due to the potential for hallucination and outdated information [12, 21, 30].
@@
Zero-shot prompting involves providing the LLMs with only the task description and the input claim, without any specific examples. This can include asking the model to directly predict the veracity label of a claim [12, 71]. Frameworks such as FACT-AUDIT [45] used zero-shot inference to evaluate the fact-checking capacity of various LLMs. They gave the lowest average accuracy scores compared to other methods, particularly when relying solely on the model’s internal knowledge. While a simple combination of self-consistency and zero-shot prompt was found to be the most effective overall strategy in a multilingual fact-checking study, this effectiveness was strongly tied to the self-consistency decoding strategy, not necessarily the zero-shot nature itself [32]. On the other hand, frameworks such as PACAR [14], which employed explicit claim decomposition and a dynamic planning mechanism, achieved strong performance in zero-shot settings and outperformed other LLM-based approaches, including few-shot and fine-tuned methods. However, relying solely on internal knowledge for fact-checking is considered unreliable and insufficient, as LLMs are prone to hallucination [12, 21, 30]. Zero-shot prompting without external access means that the model must rely on potentially inaccurate or outdated information stored during training [48]. The analysis of rationales generated through zero-shot CoT indicates unreliability for factuality analysis based on internal memorization [35]. The most significant mitigation discussed is the incorporation of external knowledge retrieval [15, 20, 21, 30, 48, 53].
The Few-Shot Prompting or ICL method involves providing the model with a limited number of examples of the task before presenting the claim to be verified [12, 14]. It utilizes the LLM’s ability to learn from examples provided directly in the prompt ("in-context learning") [20]. Few-shot demonstrations are used in methods like HiSS [21] and BiDeV [37] to guide the LLM through multi-step processes. Few-shot-CoT includes example pairs to guide the reasoning process [47, 71]. A similar approach was used to address the challenge of identifying hallucinations in natural language generation by using the pre-trained Llama 2-7B model in a zero-shot setting. Instead of fine-tuning the model, the authors used specific prompts that included the context and hypothesis sentences, which were provided to the model, to determine whether the context supported the hypothesis with a simple yes or no. The model’s answers were reprocessed to classify them and determine hallucinations or unsupported outputs. They evaluated its performance on two tracks: the model-aware method, with access to specific model checkpoints, and the model-agnostic method, without such access. To keep the process efficient, the method focused on minimal engineering and prevented additional fine-tuning or preprocessing. They claim that the model’s ability to reliably identify hallucinations could be improved even further with adjustments such as fine-tuning it for a specific domain or incorporating more data [115]. In another instance, an equivalent method is used in a prompt-based approach to few-shot learning to identify hallucinations in Swedish and English paraphrased texts using the Mistral 7B LLM. Task-specific prompts are carefully designed to include instructions and a limited number of reference examples from the training data. The model is given these prompts, the source sentence, and two hypotheses. Then, it is asked to determine which hypothesis contains hallucinated content. Without any additional fine-tuning, the method makes effective use of the model’s reasoning abilities and strategically incorporates representative samples into the prompt to help the model make decisions that will successfully detect hallucinations [116].
@@
ICL has been demonstrated to enhance the performance of open-source multimodal LLMs (MLLMs) in misinformation detection, occasionally yielding greater improvements than those achieved through prompt ensemble methods [57]. Combining ICL with RAG has been shown to improve accuracy in fact verification [53]. However, some sophisticated few-shot ICL methods like Standard Prompting, Vanilla CoT, and ReAct were surpassed by the HiSS method in news claim verification, highlighting the importance of the specific method prompted [48]. While ICL improves performance, open-source MLLMs using ICL still significantly lag behind state-of-the-art proprietary models like GPT-4V. The effectiveness can vary between models and datasets [57]. The manual prompt design for few-shot examples can be heuristic [20]. However, providing examples does not guarantee overcoming the dependence on internal knowledge if external information is not integrated [21]. Integrating ICL with RAG or structured step-by-step prompting frameworks is a key mitigation [21, 53].
@@
Other methods, such as CoT prompting, instruct the LLM to output a sequence of intermediate reasoning steps before arriving at the final answer or the veracity label [12, 21]. Zero-shot CoT prompts often include eliciting sentences like "Let us think step by step" [12, 16, 32, 47]. This encourages explicit reasoning. Variants of CoT prompting include English CoT (EN-CoT) [32], which focuses on monolingual reasoning, and CoTVP (CoT Veracity Prediction) [22], which evaluates the truthfulness of reasoning steps. However, studies show that techniques like CoT, designed to improve reasoning, do not necessarily improve the fact-checking abilities of LLMs. They can even have minimal or negative effects on success rates [32]. Models prompted with CoT may tend to align with the input text rather than verifying its factualness, especially for complex paragraphs, when relying solely on pre-trained knowledge [20]. Vanilla CoT suffers substantially from the issues of fact hallucination and omission of necessary thoughts in the reasoning process [21]. Using the LLM for factuality analysis based on its internal memorization, even with zero-shot CoT, indicates unreliability, likely caused by hallucination [12].
@@
The internal mechanism of LLMs to integrate rationales from various perspectives via CoT can be ineffective for fake news detection [12]. Integrating CoT with external knowledge retrieval, as in Search-Augmented CoT or ReAct [21], is a key approach to mitigate hallucination and thought omission [21]. Furthermore, frameworks that explicitly guide the decomposition and reasoning process, such as FactAgent, are seen as superior to CoT, which primarily acts as a prompting technique [71].
@@
When relying solely on internal knowledge, LLMs prompted with zero-shot, few-shot, or vanilla CoT struggle in fact-checking complex claims and exhibit hallucination [12, 20, 21, 30]. However, accuracy can be low, and improvements in reasoning via CoT alone do not guarantee better fact-checking performance [32]. Techniques for forcing binary "true" or "false" judgments also do not enhance overall accuracy [72].
@@ 4.4.2  Prompting Strategies with Integrated External Retrieval
A variant of CoT that interleaves reasoning traces with task‑specific actions, like querying Google Search or the Wikipedia API, allowing the LLM agent to decide whether to search or continue reasoning based on environmental observations [20, 21, 22, 48]. For instance, by accessing external knowledge, ReAct effectively mitigates hallucination failures compared to vanilla CoT and justifies its reasoning with retrieved citations, enhancing verifiability and explainability. Its performance is highly sensitive to the quality and relevance of search results, and relying solely on internal knowledge when external search fails remains a key limitation [21]. Combining ReAct’s action capabilities with more structured decomposition and step‑by‑step verification methods, such as HiSS or SELF‑CHECKER, can address thought omissions and improve overall performance [20, 21].
@@
Search‑Augmented CoT augments vanilla CoT by using the original claim as a search query to retrieve background information, which the LLM incorporates into its thought chain. This approach improves over vanilla CoT by utilizing external knowledge, but can fall short of methods like Standard Prompting or HiSS, which indicates that querying solely with the claim may output insufficiently detailed results. To mitigate this, more sophisticated query generation strategies and integration methods are needed [21]. HiSS is a few‑shot method that prompts the LLM to perform claim verification in fine‑grained steps by decomposing claims into subclaims and verifying each step‑by‑step, raising questions and optionally using web search when confidence is low. It significantly surpasses few‑shot ICL counterparts like Standard Prompting, Vanilla CoT, and ReAct in average F1-score, offering superior explainability through enhanced coverage and readability while substantially reducing hallucination and thought omission [21]. However, HiSS still struggles with integrating updated information from mixed sources and incurs high computational costs due to multiple LLM calls, with its performance remaining sensitive to prompt design [20].
@@
Frameworks like SELF‑CHECKER [20], BiDeV [37], RAGAR [22], and PACAR [14] integrate prompting and RAG by decomposing fact‑checking into subtasks (e.g., claim detection, retrieval, sentence selection, verdict prediction) and using prompts often with few‑shot examples to generate search queries, select evidence, and perform step‑by‑step verification while explicitly incorporating retrieved documents. Incorporating external knowledge via RAG significantly boosts accuracy over internal‑only approaches: SELF‑CHECKER [20], BiDeV [37], RAGAR [22], and PACAR [14] all show substantial gains, with specialized modules such as the Claim Atomizer and Fact‑Check‑Then‑RAG further enhancing predictive power and explanation quality.
@@
RAG‑based methods can be hampered by overwhelming context windows [48], outdated or variable search results [20, 22], high computational costs, prompt sensitivity, and manual prompt design [20], and they may still miss fine‑grained details even when grounded [30]. To mitigate these issues, current strategies include using IR functions (e.g., BM25) to distill relevant content [30, 48], using multi‑agent [37, 75], explicit decomposition and filtering [14, 37, 42], feedback loops [51], and refined RAG processes [30, 42]. Reliance solely on an LLM’s internal knowledge, whether via zero‑shot, few‑shot, or vanilla CoT prompting, is unreliable for fact‑checking and prone to substantial hallucination [12, 21, 30], as zero‑shot CoT and vanilla CoT often succumb to memorization pitfalls [12, 16]. Mitigating hallucination primarily involves incorporating external knowledge through ReAct, Search‑Augmented CoT, HiSS, and other RAG frameworks [21, 30, 53].
@@
In conclusion, while basic prompting strategies like zero-shot, few-shot, and vanilla CoT offer foundational ways to interact with LLMs for fact-checking, their accuracy and reliability are severely limited by the reliance on potentially flawed internal knowledge [12, 21, 30]. The most effective approaches, as highlighted by the sources, involve prompting strategies that explicitly integrate external knowledge retrieval through frameworks such as ReAct, HiSS, SELF-CHECKER, BiDeV, RAGAR, and PACAR [20, 21, 37, 48, 53]. These methods use prompts to guide LLM through processes that involve external data access, significantly improving accuracy and mitigating hallucination by grounding the model’s responses in evidence [20, 21, 53].